OpenStudy (anonymous):

i = $125 r = 6% t = 1 What is p? $56.25 $2,027.08 $2,083.33

OpenStudy (anonymous):

Do you have anymore information?

OpenStudy (anonymous):

interest=principal*rate*time

OpenStudy (anonymous):

So far you have 125= x*.06*1. We need to solve for x

OpenStudy (anonymous):

It makes it easier because we have a one at the end. So now we work backwards. 125/.06 (Make sure to divide not multiply or your answer will be wrong. So we get 2083.33 repeating. Now we input that and solve. 125=2083.33*.06*1 125=125*1 125=125 So the answer is $2,083.33

OpenStudy (anonymous):

thanku sir

OpenStudy (anonymous):

No problem.